# Towards Precise 3D Quantum Control of a Levitated Dipolar Scatterer using Spatial Mode Decomposition

### v1

@@ Acknowledgement
This work was supported by the Australian Research Council Centre of Excellence for Engineered Quantum Systems (Grant No. CE170100009) and Lockheed Martin. C. L. acknowledges support from the Sydney Quantum Academy Postdoctoral Research Fellowship. M. K. S. acknowledges support from the Macquarie University Research Fellowship scheme (MQRF0001036) and the Australian Research Council Discovery Early Career Researcher Award (DE220101272).
Supplementary Material: Towards Precise 3D Quantum Control of a Levitated Dipolar Scatterer using Spatial Mode Decomposition
I Decomposition of the Scattered Field
We are interested in calculating the field scattered from a levitated dipolar scatterer, having been transformed by the optical trap design considered in the main text. In the Fraunhofer approximation, the scattered field is analogous to the far-field of a dipole, with an appropriate phase-shift. That is,
$\bm{E}_{\mathrm{scat}}(\bm{r},\bm{q}(t))\approx\bm{E}_{\mathrm{dip}}(\bm{r})~{}\mathrm{exp}\left[-ik\cdot\left(\frac{\bm{r}\cdot\bm{q}(t)}{|\bm{r}|}\right)\right]~{},$ (S1)
@@ I Decomposition of the Scattered Field
$\bm{E}_{\mathrm{dip}}(\bm{r})$ $=\frac{k^{2}}{4\pi\varepsilon_{0}}\frac{\mathrm{exp}\Big{[}ik|\bm{r}|\Big{]}}{|\bm{r}|}\Big{[}\bm{r}\times\bm{p}(\bm{r})\Big{]}\times\bm{r}~{}.$ (S2)
Here, $\bm{q}(t)$ refers to the nanoparticle’s real-time displacement from the focal point of the parabolic mirror, and $\bm{p}(\bm{r})$ to the induced dipole moment, which we assume is aligned parallel to the polarization axis of the trapping beam. Eq. (S1) does not describe the electric field observed at the input to the space-division de-multiplexing photonic chip, however. Instead, the scattered field is transformed by a series of optical elements, resulting in modifications to both the spatially-varying phase and amplitude. In particular, for the optical trap design considered here, the back-scattered component of the field is first collimated by a gold-coated, high-numerical aperture (NA) parabolic mirror. We write the effective NA of this device as being of the form
$\mathrm{NA}$ $=1-\mathrm{cos}\left(\mathrm{arctan}\left[\frac{r_{\mathrm{m}}}{f_{\mathrm{m}}-\frac{r_{\mathrm{m}}^{2}}{4f_{\mathrm{m}}}}\right]\right)~{},$ (S3)
where $f_{\mathrm{m}}$ and $r_{\mathrm{m}}$ are the focal length (1 mm) and radius (2 mm) of the parabolic mirror, respectively [16]. The collimated field then propagates to some collection lens of focal length $f_{\mathrm{cl}}$ (25 mm) and is subsequently focused in the input plane of the photonic chip. Following derivations conducted in the context of confocal microscopy, we opt to calculate the transformed field using angular spectrum theory [26]. For the setup considered here, we find a time-independent component of the electric field, in the focal plane of the collection lens, of the form
$\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}(t))$ $=-\frac{ik^{3}f_{\mathrm{cl}}}{16\pi^{3}\varepsilon_{0}}\int^{\theta_{\mathrm{cl}}^{\mathrm{max}}}_{0}\int^{2\pi}_{0}\frac{[1+\mathrm{cos}\left(\theta_{\mathrm{m}}\right)]}{2f_{\mathrm{m}}}\sqrt{\mathrm{cos}\left(\theta_{\mathrm{cl}}\right)}\left[(\bm{p}(\bm{r})\cdot\bm{e}_{\mathrm{m}}^{\parallel})\cdot\bm{e}_{\mathrm{cl}}^{\parallel}-(\bm{p}(\bm{r})\cdot\bm{e}_{\perp})\cdot\bm{e}_{\perp}\right]$
×$\hskip 99.58464pt\times\mathrm{exp}\Big{[}ik(x~{}\mathrm{sin}(\theta_{\mathrm{m}})\mathrm{cos}(\varphi)+y~{}\mathrm{sin}(\theta_{\mathrm{m}})\mathrm{sin}(\varphi)-(\hat{\bm{e}}_{q}\cdot\bm{s}_{\mathrm{m}})q(t)\Big{]}\mathrm{sin}(\theta_{\mathrm{cl}})~{}\mathrm{d}\theta_{\mathrm{cl}}\mathrm{d}\varphi~{},$ (S4)
@@
$\bm{e}_{\perp}$ $=\Big{[}-\mathrm{sin}(\varphi),~{}\mathrm{cos}(\varphi),~{}0\Big{]}$ (S5)
$\bm{e}_{\mathrm{m}}^{\parallel}$ $=\Big{[}\mathrm{cos}(\theta_{\mathrm{m}})\mathrm{cos}(\varphi),~{}\mathrm{cos}(\theta_{\mathrm{m}})\mathrm{sin}(\varphi),~{}\mathrm{sin}(\theta_{\mathrm{m}})\Big{]}$ (S6)
$\bm{e}_{\mathrm{cl}}^{\parallel}$ $=\Big{[}-\mathrm{cos}(\theta_{\mathrm{cl}})\mathrm{cos}(\varphi),~{}-\mathrm{cos}(\theta_{\mathrm{cl}})\mathrm{sin}(\varphi),~{}\mathrm{sin}(\theta_{\mathrm{cl}})\Big{]}$ (S7)
$\bm{s}_{\mathrm{m}}$ $=\Big{[}-\mathrm{sin}(\theta_{\mathrm{cl}})\mathrm{cos}(\varphi),~{}-\mathrm{sin}(\theta_{\mathrm{cl}})\mathrm{sin}(\varphi),~{}\mathrm{cos}(\theta_{\mathrm{cl}})\Big{]}~{}.$ (S8)
@@
$\theta_{\mathrm{m}}$ $=2~{}\mathrm{arctan}\left(\frac{M~{}\mathrm{sin}(\theta_{\mathrm{cl}})}{2}\right)$ (S9)
$\theta_{\mathrm{cl}}^{\mathrm{max}}$ $=\mathrm{arctan}\left(\frac{2r_{\mathrm{m}}}{f_{\mathrm{cl}}}\right)~{},$ (S10)
where $M=25$ and $\theta_{\mathrm{cl}}^{\mathrm{max}}\approx 0.16$ radians for the particular experimental setup considered here. Taylor expanding the exponential term in equation (S4) about $q(t)=0$, we find that we can write the scattered field as
$\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}(t))$ $=\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}=\bm{0})+\sum_{n>0}\frac{q(t)^{n}}{w_{0}^{n}}\frac{1}{n!}\left[\frac{\partial^{n}\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}(t))}{\partial q(t)^{n}}\right]_{q(t)=0}$
$=\bm{E}_{0}(x,y)+\sum_{n>0}\frac{q(t)^{n}}{w_{0}^{n}}\bm{E}_{\mathrm{n}}(x,y,\bm{\hat{e}}_{q})~{},$ (S11)
where $w_{0}$ refers to the beam waist at the focal point of the parabolic mirror, and has been introduced to normalize $q(t)$. Upon inspection, we find that this is equivalent to decomposing the total field into elastically and in-elastically scattered contributions. Restricting ourselves to first order, and generalising to the case where each translational degree of freedom is de-coupled from one another, this becomes
$\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}(t))$ $=\bm{E}_{0}(x,y)+\sum_{q=x,y,z}\frac{q(t)}{w_{0}}\bm{E}^{(q)}(x,y)~{}.$ (S12)
II Calculating the Coupling Efficiencies
To calculate the coupling of each scattered field component into the space-division de-multiplexing photonic chip (spatial mode sorter), we must first define the guided modes. The input to the chip is an embedded few-mode, graded-index optical fiber (FMF) designed for operation at $\lambda=1550$ nm, with a core radius ($r_{\mathrm{fib}}$) of approximately 4 µm. We treat the guided modes as linearly-polarized (LP) modes, whose electric field distributions are fully described by the family of complex amplitude functions,
$\mathrm{LP}_{\ell,m}(x,y,z)=\begin{cases}J_{\ell}(\frac{ur}{r_{\mathrm{fib}}})\mathrm{cos}\left(\ell\phi\right)~{}\mathrm{exp}\big{[}-i\beta_{\ell,m}z\big{]}&,~{}r\leq r_{\mathrm{fib}}\\ \frac{J_{\ell}(u)}{K_{\ell}(w)}K_{\ell}(\frac{wr}{r_{\mathrm{fib}}})\mathrm{cos}\left(\ell\phi\right)~{}\mathrm{exp}\big{[}-i\beta_{\ell,m}z\big{]}&,~{}r\geq r_{\mathrm{fib}}\end{cases}$ (S13)
@@ II Calculating the Coupling Efficiencies
$u$ $=r_{\mathrm{fib}}\sqrt{n_{1}^{2}k^{2}-\beta_{\ell,m}^{2}}$ (S14)
$w$ $=r_{\mathrm{fib}}\sqrt{\beta_{\ell,m}^{2}-n_{2}^{2}k^{2}}~{},$ (S15)
@@
in terms of the propagation constant, $\beta_{\ell,m}$. Further, $J_{\ell}(\bullet)$ and $K_{\ell}(\bullet)$ refer to Bessel functions of the first kind, and modified Bessel functions of the second kind, respectively. In our particular experimental setup, the spatial profile of each LP mode is rotated clockwise through an angle of approximately $\theta=16^{\mathrm{o}}$ from the laboratory frame of reference. Additionally, the chip is inclined by approximately $\phi=5^{\mathrm{o}}$ to the propagation axis of the trapping beam. This introduces a weak longitudinally-polarised component to each LP mode, which plays a significant role in our detection technique. To this end, we define the set of electric fields,
@@
$\bm{E}_{\mathrm{fib}}^{\ell,m}(x,y)$ $=\mathrm{LP}_{\ell,m}\Big{(}x~{}\mathrm{cos}\left(\theta\right)-y~{}\mathrm{sin}\left(\theta\right),x~{}\mathrm{sin}\left(\theta\right)+y~{}\mathrm{cos}\left(\theta\right),0\Big{)}\left(\begin{smallmatrix}1\\ \mathrm{cos}\left(\phi\right)\\ \mathrm{sin}\left(\phi\right)\end{smallmatrix}\right)~{},$ (S16)
which form a complete basis, being mutually orthonormal with respect to the spatial-overlap integral. To calculate the coupling of the in-elastically scattered field $\bm{E}^{(q)}(x,y)$ into any given LP mode, we calculate the quantity
@@
$\eta_{q}^{\ell,m}$ $=\frac{\left|\iint_{\mathrm{FMF}}\big{[}\bm{E}^{(q)}(x,y)\big{]}^{*}\cdot\bm{E}_{\mathrm{fib}}^{\ell,m}(x,y)~{}\mathrm{d}x\mathrm{d}y\right|}{\sqrt{\iint_{\mathrm{FMF}}\big{|}\bm{E}^{(q)}(x,y)\big{|}^{2}~{}\mathrm{d}x\mathrm{d}y\times\iint_{\mathrm{FMF}}\big{|}\bm{E}_{\mathrm{fib}}^{\ell,m}(x,y)\big{|}^{2}~{}\mathrm{d}x\mathrm{d}y}}~{},$ (S17)
where integration is carried out over the end-face of the few-mode input fiber (FMF).
III Deriving the Measurement Imprecision
We must realistically consider the coupling of the total scattered field into the spatial mode sorter, not just the in-elastically scattered term $\bm{E}^{(q)}(x,y)$. Assuming no internal losses, we expect that a detector measures an output power in the LPℓ,m channel of
$P^{\ell,m}_{q}(t)$ $=\sqrt{\frac{\varepsilon_{0}}{\mu_{0}}}\iint_{\mathrm{FMF}}\left|\eta^{\ell,m}_{0}\bm{E}_{0}(x,y)+\frac{q(t)}{w_{0}}\eta^{\ell,m}_{q}\bm{E}^{(q)}(x,y)+\eta^{\ell,m}_{\mathrm{ref}}\bm{E}_{\mathrm{ref}}(x,y)\right|^{2}~{}\mathrm{d}x\mathrm{d}y$
$=\sqrt{\frac{\varepsilon_{0}}{\mu_{0}}}\iint_{\mathrm{FMF}}(\eta^{\ell,m}_{\mathrm{ref}})^{2}\left|\bm{E}_{\mathrm{ref}}(x,y)\right|^{2}+(\eta^{\ell,m}_{0})^{2}\left|\bm{E}_{0}(x,y)\right|^{2}+2\eta^{\ell,m}_{\mathrm{ref}}\eta^{\ell,m}_{0}\left|\bm{E}_{\mathrm{ref}}^{*}(x,y)\bm{E}_{0}(x,y)\right|$
$\hskip 72.83907pt+2\frac{q(t)}{w_{0}}\eta^{\ell,m}_{\mathrm{ref}}\eta^{\ell,m}_{q}\left|\bm{E}_{\mathrm{ref}}^{*}(x,y)\bm{E}^{(q)}(x,y)\right|$
$\hskip 72.83907pt+2\frac{q(t)}{w_{0}}\eta^{\ell,m}_{0}\eta^{\ell,m}_{q}\left|\bm{E}_{0}^{*}(x,y)\bm{E}^{(q)}(x,y)\right|~{}\mathrm{d}x\mathrm{d}y~{},$ (S18)

### v2

@@ Acknowledgement
Authors acknowledge fruitful discussions with Thiago Guerreiro, Michael Steel and Massimiliano Rossi. This work was supported by the Australian Research Council Centre of Excellence for Engineered Quantum Systems (Grant No. CE170100009) and Lockheed Martin. C. L. acknowledges support from the Sydney Quantum Academy Postdoctoral Research Fellowship. M. K. S. acknowledges support from the Australian Research Council Discovery Early Career Researcher Award (DE220101272).
Supplementary Material: Three-dimensional quantum-enabling displacement sensing of a levitated nanoparticle
Appendix A Decomposition of the scattered field
In this Appendix we provide the explicit expressions for the field scattered from a levitated dipolar scatterer, having been transformed by the parabolic mirror and collection lens composing the optical trap. In the Fraunhofer approximation, the scattered field is analogous to the far-field of an electric dipole, with an appropriate phase-shift:
$\bm{E}_{\mathrm{scat}}(\omega,\bm{r}~{};~{}\bm{q}(t))\approx\bm{E}_{\mathrm{dip}}(\omega,\bm{r})~{}\mathrm{exp}\left[-ik\left(\frac{\bm{r}\cdot\bm{q}(t)}{|\bm{r}|}\right)\right]~{},$ (4)
@@ Appendix A Decomposition of the scattered field
$\bm{E}_{\mathrm{dip}}(\omega,\bm{r})$ $=\frac{k^{2}}{4\pi\varepsilon_{0}}\frac{e^{ik|\bm{r}|}}{|\bm{r}|}\big{[}\bm{r}\times\bm{p}(\omega,\bm{r})\big{]}\times\bm{r}~{}.$ (5)
Here, $\bm{q}(t)$ refers to the nanoparticle’s real-time displacement from the focal point of the parabolic mirror, and $\bm{p}(\omega,\bm{r})$ to the induced dipole moment, which we assume is aligned parallel to the polarization axis of the trapping beam. Moving forward, we drop the frequency dependence of these terms, and choose to focus on the case of a monochromatic trapping beam. We note that Eq. (4) does not describe the electric field observed at the input to the space-division demultiplexing photonic chip, however. Instead, the scattered field is transformed by a series of optical elements, resulting in modifications to both the spatially-varying phase and amplitude. In particular, for the optical trap design considered here, the back-scattered component of the field is first collimated by a gold-coated, high-numerical aperture (NA) parabolic mirror. We write the effective NA of this device as being of the form
$\mathrm{NA}$ $=1-\mathrm{cos}\left(\mathrm{arctan}\left[\frac{r_{\mathrm{m}}}{f_{\mathrm{m}}-\frac{r_{\mathrm{m}}^{2}}{4f_{\mathrm{m}}}}\right]\right)~{},$ (6)
where $f_{\mathrm{m}}$ and $r_{\mathrm{m}}$ are the focal length (1 mm) and radius (2 mm) of the parabolic mirror, respectively [21]. The collimated field then propagates to some collection lens of focal length $f_{\mathrm{cl}}$ (25 mm) and is subsequently focused in the input plane of the photonic chip. A schematic of this setup is shown in Fig. 4 below.
Figure 4: Geometry of the parabolic mirror being modelled, with key features labelled. A nanoparticle trapped at the diffraction-limited focal point of the parabolic mirror scatters a portion of the incident light, which is then coupled into the few-mode fiber input of a space-division demultiplexing photonic chip.
This problem, of imaging the light scattered from a dipolar source, has been studied extensively in the context of confocal microscopy [36], where it is typically addressed using angular spectrum theory. For the setup considered here, we find a time-independent component of the electric field, in the focal plane of the collection lens, of the form
$\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}(t))$ $=-\frac{ik^{3}f_{\mathrm{cl}}}{16\pi^{3}\varepsilon_{0}}\int^{\theta_{\mathrm{cl}}^{\mathrm{max}}}_{0}\int^{2\pi}_{0}\frac{[1+\mathrm{cos}\left(\theta_{\mathrm{m}}\right)]}{2f_{\mathrm{m}}}\sqrt{\mathrm{cos}\left(\theta_{\mathrm{cl}}\right)}\left[(\bm{p}\cdot\bm{e}_{\mathrm{m}}^{\parallel})\cdot\bm{e}_{\mathrm{cl}}^{\parallel}-(\bm{p}\cdot\bm{e}_{\perp})\cdot\bm{e}_{\perp}\right]$
×$\hskip 99.58464pt\times\mathrm{exp}\Big{[}-ik(x~{}\mathrm{sin}(\theta_{\mathrm{cl}})\mathrm{cos}(\varphi)+y~{}\mathrm{sin}(\theta_{\mathrm{cl}})\mathrm{sin}(\varphi)+(\hat{\bm{e}}_{q}\cdot\bm{s}_{\mathrm{m}})q(t))\Big{]}\mathrm{sin}(\theta_{\mathrm{cl}})~{}\mathrm{d}\theta_{\mathrm{cl}}\mathrm{d}\varphi~{},$ (7)
@@
$\bm{e}_{\perp}$ $=\Big{[}-\mathrm{sin}(\varphi),~{}\mathrm{cos}(\varphi),~{}0\Big{]}$ (8)
$\bm{e}_{\mathrm{m}}^{\parallel}$ $=\Big{[}\mathrm{cos}(\theta_{\mathrm{m}})\mathrm{cos}(\varphi),~{}\mathrm{cos}(\theta_{\mathrm{m}})\mathrm{sin}(\varphi),~{}\mathrm{sin}(\theta_{\mathrm{m}})\Big{]}$ (9)
$\bm{e}_{\mathrm{cl}}^{\parallel}$ $=\Big{[}-\mathrm{cos}(\theta_{\mathrm{cl}})\mathrm{cos}(\varphi),~{}-\mathrm{cos}(\theta_{\mathrm{cl}})\mathrm{sin}(\varphi),~{}-\mathrm{sin}(\theta_{\mathrm{cl}})\Big{]}$ (10)
$\bm{s}_{\mathrm{m}}$ $=\Big{[}\mathrm{sin}(\theta_{\mathrm{m}})\mathrm{cos}(\varphi),~{}\mathrm{sin}(\theta_{\mathrm{m}})\mathrm{sin}(\varphi),~{}-\mathrm{cos}(\theta_{\mathrm{m}})\Big{]}~{}.$ (11)
@@
$\theta_{\mathrm{m}}$ $=2~{}\mathrm{arctan}\left(\frac{M~{}\mathrm{sin}(\theta_{\mathrm{cl}})}{2}\right)$ (12)
$\theta_{\mathrm{cl}}^{\mathrm{max}}$ $=\mathrm{arctan}\left(\frac{2r_{\mathrm{m}}}{f_{\mathrm{cl}}}\right)~{},$ (13)
where $M=25$ and $\theta_{\mathrm{cl}}^{\mathrm{max}}\approx 0.16$ radians for the particular experimental setup considered here. Taylor expanding the exponential term in Eq. (7) to first-order about $\bm{q}=\bm{0}$, we find that we can write the scattered field as
$\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}(t))$ $=\bm{E}_{\mathrm{scat}}(x,y~{};~{}\bm{q}=\bm{0})~{}+\stackrel{{\scriptstyle\leftrightarrow}}{{\bm{J}}}_{\mathrm{scat}}(x,y~{};~{}\bm{q}=\bm{0})\bm{q}(t)^{T}~{},$ (14)
where we have introduced the Jacobian tensor
$\stackrel{{\scriptstyle\leftrightarrow}}{{\bm{J}}}_{\mathrm{scat}}(x,y~{};~{}\bm{q}=\bm{0})$ $=\begin{bmatrix}\nabla^{T}\bm{E}_{\mathrm{scat}}(x,y~{};~{}\bm{q}=\bm{0})\cdot\hat{\bm{e}}_{x}\\ \nabla^{T}\bm{E}_{\mathrm{scat}}(x,y~{};~{}\bm{q}=\bm{0})\cdot\hat{\bm{e}}_{y}\\ \nabla^{T}\bm{E}_{\mathrm{scat}}(x,y~{};~{}\bm{q}=\bm{0})\cdot\hat{\bm{e}}_{z}\end{bmatrix}~{},$ (15)
for
$\nabla^{T}f=\left[\frac{\partial f}{\partial(\bm{q}\cdot\hat{\bm{e}}_{x})},\frac{\partial f}{\partial(\bm{q}\cdot\hat{\bm{e}}_{y})},\frac{\partial f}{\partial(\bm{q}\cdot\hat{\bm{e}}_{z})}\right]~{}.$ (16)
For the simplest case where the nanoparticle oscillates about a single spatial axis, such that $\bm{q}(t)=q(t)\hat{\bm{e}}_{q}$ for some principle axis $q=x,y$ or $z$, Eq. (14) reduces to
$\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}(t))$ $=\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}=\bm{0})+q(t)\frac{\partial\bm{E}_{\mathrm{scat}}(x,y~{};~{}\bm{q}=\bm{0})}{\partial(\bm{q}\cdot\hat{\bm{e}}_{q})}$
$=\bm{E}_{0}(x,y)+\frac{q(t)}{w_{0}}\bm{E}^{(q)}(x,y)~{},$ (17)
where $w_{0}$ refers to the beam waist at the focal point of the parabolic mirror, and has been introduced to normalize $q(t)$. Upon inspection, we find that this is equivalent to decomposing the total field into elastically and inelastically scattered contributions.
Appendix B Calculating the coupling efficiencies
To calculate the coupling of each scattered field component into the space-division demultiplexing photonic chip (spatial mode sorter), we need to formulate the guided modes. The input to the chip is an embedded few-mode, graded-index optical fiber (FMF) designed for operation at $\lambda=1550$ nm, with a core radius ($r_{\mathrm{fib}}$) of approximately 4 µm. Due to the low refractive index contrast between the refractive indices of the core ($n_{1}$) and cladding ($n_{2}$) we treat the guided modes as linearly-polarized (LP) modes, whose electric field distributions are fully described by the family of complex amplitude functions
$\mathrm{LP}_{\ell,m}(\bm{r})=\begin{cases}J_{\ell}\Big{(}\frac{u\sqrt{x^{2}+y^{2}}}{r_{\mathrm{fib}}}\Big{)}\mathrm{cos}\left(\ell\phi\right)~{}\mathrm{exp}\big{[}-i\beta_{\ell,m}z\big{]}&,~{}\sqrt{x^{2}+y^{2}}\leq r_{\mathrm{fib}}\\ \frac{J_{\ell}(u)}{K_{\ell}(w)}K_{\ell}\Big{(}\frac{w\sqrt{x^{2}+y^{2}}}{r_{\mathrm{fib}}}\Big{)}\mathrm{cos}\left(\ell\phi\right)~{}\mathrm{exp}\big{[}-i\beta_{\ell,m}z\big{]}&,~{}\sqrt{x^{2}+y^{2}}\geq r_{\mathrm{fib}}\end{cases}$ (18)
@@ Appendix B Calculating the coupling efficiencies
$u$ $=r_{\mathrm{fib}}\sqrt{n_{1}^{2}k^{2}-\beta_{\ell,m}^{2}}$ (19)
$w$ $=r_{\mathrm{fib}}\sqrt{\beta_{\ell,m}^{2}-n_{2}^{2}k^{2}}~{},$ (20)
@@
in terms of the propagation constant, $\beta_{\ell,m}$. Here, $J_{\ell}(\bullet)$ and $K_{\ell}(\bullet)$ refer to Bessel functions of the first kind, and modified Bessel functions of the second kind, respectively. In our particular setup, the spatial profile of each LP mode is rotated anti-clockwise through an angle of approximately $\theta=14^{\mathrm{o}}$ from the laboratory frame of reference. To account for this, we transform the vector field of each LP mode according to
$\textbf{LP}_{\ell,m}(\bm{r})$ $=R(\theta)~{}\mathrm{LP}_{\ell,m}(\bm{r}^{\prime})~{}\hat{\bm{e}}_{x}~{},$ (21)
where $\bm{r}^{\prime}=R^{T}(\theta)~{}\bm{r}$ and we have assumed that we are concerned with only those LP modes polarised along the $\hat{\bm{e}}_{x}$ axis. To perform this calculation, we have made use of the usual rotation matrix
$R(\theta)$ $=\Bigg{[}\begin{smallmatrix}\mathrm{cos}\left(\theta\right)&\mathrm{sin}\left(\theta\right)&0\\ -\mathrm{sin}\left(\theta\right)&\mathrm{cos}\left(\theta\right)&0\\ 0&0&1\end{smallmatrix}\Bigg{]}~{}.$ (22)
Additionally, to accommodate for the angled end-face of the FMF, the photonic chip is inclined by approximately $\phi=5^{\mathrm{o}}$ to the propagation axis. This introduces a weak longitudinally-polarised component to each LP mode. Hence, repeating the calculations above for a secondary rotation matrix
$R(\phi)$ $=\Bigg{[}\begin{smallmatrix}1&0&0\\ 0&\mathrm{cos}\left(\phi\right)&\mathrm{sin}\left(\phi\right)\\ 0&-\mathrm{sin}\left(\phi\right)&\mathrm{cos}\left(\phi\right)\end{smallmatrix}\Bigg{]}~{},$ (23)
we may define a set of electric fields
@@
$\textbf{LP}_{\ell,m}(\bm{r})$ $=\mathrm{LP}_{\ell,m}(\bm{r}^{\prime\prime})\Bigg{[}\begin{smallmatrix}\mathrm{cos}\left(\theta\right)\\ -\mathrm{sin}\left(\theta\right)\mathrm{cos}\left(\phi\right)\\ \mathrm{sin}\left(\theta\right)\mathrm{sin}\left(\phi\right)\end{smallmatrix}\Bigg{]}~{},$ (24)
in terms of the modified coordinates
$\bm{r}^{\prime\prime}$ $=\Bigg{[}\begin{smallmatrix}x~{}\mathrm{cos}\left(\theta\right)-y~{}\mathrm{sin}\left(\theta\right)\\ \mathrm{cos}\left(\phi\right)[x~{}\mathrm{sin}\left(\theta\right)+y~{}\mathrm{cos}\left(\theta\right)]-z~{}\mathrm{sin}\left(\phi\right)\\ \mathrm{sin}\left(\phi\right)[x~{}\mathrm{sin}\left(\theta\right)+y~{}\mathrm{cos}\left(\theta\right)]+z~{}\mathrm{cos}\left(\phi\right)\end{smallmatrix}\Bigg{]}~{}.$ (25)
Importantly, these fields remain mutually orthogonal with respect to the normalized inner-product
$\left|\iint_{-\infty}^{\infty}\textbf{LP}_{\ell_{1},m_{1}}^{*}(x,y)\cdot\textbf{LP}_{\ell_{2},m_{2}}(x,y)~{}\mathrm{d}x\mathrm{d}y\right|$ $=0~{},$ (26)
for all $(\ell_{1},m_{1})\neq(\ell_{2},m_{2})$. To calculate the coupling of the inelastically scattered field $\bm{E}^{(q)}(x,y)$ into any given LP mode, we calculate the quantity
@@
$\eta_{q}^{\ell,m}$ $=\frac{\left|\iint_{\mathrm{FMF}}\big{[}\bm{E}^{(q)}(x,y)\big{]}^{*}\cdot\textbf{LP}_{\ell,m}(x,y)~{}\mathrm{d}x\mathrm{d}y\right|}{\sqrt{\iint_{\mathrm{FMF}}\big{|}\bm{E}^{(q)}(x,y)\big{|}^{2}~{}\mathrm{d}x\mathrm{d}y\times\iint_{\mathrm{FMF}}\big{|}\textbf{LP}_{\ell,m}(x,y)\big{|}^{2}~{}\mathrm{d}x\mathrm{d}y}}~{},$ (27)
where integration is carried out over the input face of the few-mode input fiber (FMF). Finally, we note that Eq. (27) is an approximate expression for the coupling coefficient, derived under the assumption that the coupled modes are exclusively transverse and spatially polarisation-matched [37], and that reflection from the face of the FMF is negligible. In this demonstration both conditions are approximately fulfilled, owing to the natural transverse characteristics of the LP modes, and the small contrast between air and fiber material ($n=1.4679$). Exact analysis of the coupling, employing numerical modelling of the coupling between the scattered fields and an optimised set of fiber modes, will be published elsewhere.
Appendix C Deriving the measurement imprecision
We must realistically consider the coupling of the total scattered field into the spatial mode sorter, not just the inelastically scattered term $\bm{E}^{(q)}(x,y)$. First, however, let us consider the form of the total electric field incident on the input plane of the spatial mode sorter. For the case where the nanoparticle oscillates only along the $\hat{\bm{e}}_{q}$ axis, this field is of the form
$\bm{E}_{\mathrm{inc}}(x,y)$ $=\bm{E}_{\mathrm{ref}}(x,y)+\bm{E}_{0}(x,y)+\frac{q(t)}{w_{0}}\bm{E}^{(q)}(x,y)~{}.$ (28)
Moving forward, we choose to represent the electric field in terms of the complex normalised modes ${\hat{\bm{E}}(x,y)=\bm{E}(x,y)~{}/~{}(\iint_{\mathrm{FMF}}|\bm{E}(x,y)|^{2}~{}\mathrm{d}x\mathrm{d}y)^{1/2}}$ and corresponding real-valued amplitudes $\alpha$. To this end, we write
$\bm{E}_{\mathrm{inc}}(x,y)$ $=\hat{\bm{E}}_{\mathrm{ref}}(x,y)\alpha_{\mathrm{ref}}+\hat{\bm{E}}_{0}(x,y)\alpha_{0}+\frac{q(t)}{w_{0}}\hat{\bm{E}}^{(q)}(x,y)\alpha_{q}~{}.$ (29)
The electric field excited in the LPℓ,m channel of the spatial mode sorter is then
$\textbf{LP}_{\ell,m}(x,y)$ $=\hat{\textbf{LP}}_{\ell,m}(x,y)\left[\eta^{\ell,m}_{\mathrm{ref}}\alpha_{\mathrm{ref}}+\eta^{\ell,m}_{0}\alpha_{0}+\frac{q(t)}{w_{0}}\eta^{\ell,m}_{q}\alpha_{q}\right]~{},$ (30)
where $\eta^{\ell,m}_{\mathrm{ref}}$ and $\eta^{\ell,m}_{0}$ are defined analogously to Eq. (27) except for the LO and elastically scattered fields, respectively. Neglecting any internal losses, we expect that the output power in the LPℓ,m channel measured by the detector can therefore be estimated as
$P^{\ell,m}_{q}(t)$ $=\sqrt{\frac{\varepsilon_{0}}{\mu_{0}}}\left|\eta^{\ell,m}_{0}\alpha_{0}+\frac{q(t)}{w_{0}}\eta^{\ell,m}_{q}\alpha_{q}+\eta^{\ell,m}_{\mathrm{ref}}\alpha_{\mathrm{ref}}\right|^{2}$
$=\sqrt{\frac{\varepsilon_{0}}{\mu_{0}}}\Bigg{(}(\eta^{\ell,m}_{\mathrm{ref}})^{2}\alpha_{\mathrm{ref}}^{2}+(\eta^{\ell,m}_{0})^{2}\alpha_{0}^{2}+2\eta^{\ell,m}_{\mathrm{ref}}\eta^{\ell,m}_{0}\alpha_{\mathrm{ref}}\alpha_{0}+2\frac{q(t)}{w_{0}}\eta^{\ell,m}_{\mathrm{ref}}\eta^{\ell,m}_{q}\alpha_{\mathrm{ref}}\alpha_{q}+2\frac{q(t)}{w_{0}}\eta^{\ell,m}_{0}\eta^{\ell,m}_{q}\alpha_{0}\alpha_{q}\Bigg{)},$ (31)